# Benchmarking Filtered Approximate Nearest Neighbor Search Algorithms on Transformer-based Embedding Vectors

### v1

@@ 5.8. Results on arxiv-for-fanns-large
Figure 18. 5.8) Recall@10 vs. QPS plots for the three filter types (EM, R, EMIS) on the arxiv-for-fanns-large dataset.
Figure 19. 5.8) Index construction time (left), peak memory usage (middle), and index size (right) for the three filter types (EM, R, EMIS) on the arxiv-for-fanns-large dataset.
Recall vs. QPS (Figure 18)
[…]
Index Construction Time (Figure 19, left)
Index construction on the 27×27$\times$27 × larger dataset increased the runtime of most methods by roughly an order of magnitude. Since the large dataset was processed with 9×9$\times$9 × more threads, this suggests that index construction time grows faster than linearly with dataset size or that speedup from additional threads is sublinear. CAPS remains the fastest method, while ACORN continues to lag behind most others. UNG has by far the longest construction time, which may be due to UNG failing to reach the target recall of 0.950.950.95, causing the parameter search to shift toward more expensive configurations.
Peak Memory Usage (Figure 19, middle)
Memory usage of all methods scales roughly linearly with dataset size. Relative differences during index construction and query execution remain similar to those on the medium-scale dataset.
Index Size (Figure 19, right)
UNG stands out by compressing the index to about one fourth of the dataset size without requiring access to the original vectors at query time. As in the medium-scale setting, CAPS and NHQ (kgraph) produce compact indexes of just a few megabytes but require access to the original vectors at query time. For the remaining methods, the index size is approximately equal to the dataset size.
@@ 6. Related Work
In the first part of our work, we survey the current state of the FANNS field. While we are not aware of any existing surveys specifically focused on FANNS, several works review the related ANNS domain, such as the surveys by Echihabi et al. (Echihabi et al., 2021) and Han et al. (Han et al., 2023). Other surveys explore prominent application areas of ANNS and FANNS, most notably in the context of retrieval-augmented generation (RAG) for LLMs (Gao et al., 2023) and reasoning language models (RLMs) with integrated RAG capabilities (Besta et al., 2025).
The second part of our paper introduces a novel dataset featuring transformer-based embedding vectors, designed to reflect emerging RAG workloads. This type of data is not represented in the existing landscape of ANNS and FANNS datasets (Laurent Amsaleg, 2025; Pennington et al., 2025; Dmitry Baranchuk, 2025; Simhadri et al., 2025). Our embeddings are generated using the stella_en_400M_v5 model (NovaSearch, 2025; Zhang et al., 2024), a widely used embedding model in LLM research (Besta et al., 2024; Kim et al., 2025). In contrast to our arxiv-for-fanns dataset, end-to-end RAG benchmarks such as the Massive Text Embedding Benchmark (Muennighoff et al., 2022) contain text passages rather than embedding vectors, making them unsuitable for direct evaluation of FANNS methods.
Finally, we benchmark a range of existing FANNS methods (Patel et al., 2024; Gupta et al., 2023; Wang et al., 2024; Cai et al., 2024; Gollapudi et al., 2023; Zuo et al., 2024) on our dataset. While each of these methods includes evaluation results in its respective publication, there is no comprehensive benchmark of FANNS methods on modern, transformer-based embeddings. In the broader ANNS field, several large-scale benchmarks (Aumüller et al., 2020; Aumueller et al., 2025; Wang et al., 2021c) have been established, providing valuable comparisons across methods and datasets.
@@ 7. Conclusion
The rapid progress of embedding models for various modalities has intensified the need for fast and accurate FANNS methods. In response, many approaches supporting EM, R, and EMIS filtering have emerged recently. To structure this evolving field, we present a taxonomy and survey of modern FANNS methods. We identify a key gap: the lack of open datasets with transformer-based embeddings enriched with real-world attributes. To address this, we release the arxiv-for-fanns dataset with over 2.7 million 4096-dimensional embedding vectors and 11 real-world attributes. We benchmark several FANNS methods on this dataset and analyze their performance under this challenging workload. No single method excels universally. ACORN is robust and flexible but often outperformed by specialized methods. SeRF performs well for range filters on ordered attributes but lacks categorical filtering support. Filtered-DiskANN and UNG perform well on medium-scale data but fail to scale, highlighting the difficulty of handling high-dimensional transformer-based embeddings. We conclude that the large dimensionality of transformer-based vectors poses a major challenge for FANNS methods. Moreover, we find that efficient parameter tuning remains one of the most pressing open problems in this area.

### v2

@@ 5.7.2. Results
Figure 29. 5.8) Recall@k vs. QPS plots for varying values of $k$ on the arxiv-for-fanns-medium dataset with EMIS filtering.
5.8. Analysis for Different Values of k
In Figure 27, we show recall vs. QPS plots for EM filters on the arxiv-for-fanns-medium dataset with $k\in\{20,40,60,80,100\}$. Figures 28 and 29 present the same study for R and EMIS filters, respectively. Index construction time, memory usage, and index size show almost no variation with $k$ and are therefore omitted.
Observation [Uncaptioned image]
The relative performance of different FANNS methods remains consistent across different values of $k$.
Figure 30. 5.9) Index construction time, peak memory usage, and index size for EM filtering with varying selectivity.
Figure 31. 5.9) Recall@10 vs. QPS plots for EM filtering on arxiv-for-fanns-medium with different query selectivities.
Figure 32. 5.9) Index construction time, peak memory usage, and index size for R filtering with varying selectivity.
Figure 33. 5.9) Recall@10 vs. QPS plots for R filtering on arxiv-for-fanns-medium with different query selectivities.
Figure 34. 5.9) Index construction time, peak memory usage, and index size for EMIS filtering with varying selectivity.
Figure 35. 5.9) Recall@10 vs. QPS plots for EMIS filtering on arxiv-for-fanns-medium with different query selectivities.
5.9. Analysis for Different Query Selectivities
To explore the impact of query selectivity on FANNS performance, we benchmark a version of our arxiv-for-fanns-medium dataset where we generate synthetic attributes in order to control the query selectivity. Figures 31, 33 and 35 show recall vs. QPS plots for EM, R, and EMIS filters with selectivities of $0.2$, $0.4$, $0.6$, and $0.8$. Figures 30, 32 and 34 report index construction time, peak memory usage, and index size for the same experiments. We observe that NHQ (nsw) with EM filters exhibits substantial difficulty with low selectivity scenarios, while UNG with EMIS filters handles low selectivity scenarios slightly better than high selectivity ones. There are also some variations in index construction time, with ACORN being slower to build for selectivity close to $0.5$ than for more extreme selectivities, and DIGRA being faster to build for lower selectivities. Memory usage and index size are unaffected by selectivity.
Observation [Uncaptioned image]
The performance and resource usage of most FANNS methods is largely consistent across query selectivities; however, some methods exhibit significant performance degradation in low-selectivity scenarios.
@@ 6. Related Work
Interest in FANNS is growing, with new methods introduced each year, an ACM SIGMOD programming contest (Contest, 2025) dedicated to FANNS, and survey and benchmark preprints appearing around the same time as ours. Lin et al. (Lin et al., 2025) classify methods in a pruning-oriented framework, analyzing the combination of vector pruning (i.e., ANNS) and scalar pruning (i.e., attribute filtering), while Shi et al. (Shi et al., 2025) benchmark selected methods on existing datasets, though, to the best of our knowledge, not on transformer-based embeddings. Our work complements these efforts by introducing a taxonomy along three dimensions: filtering approach, indexing technique, and filter types, and by contributing a transformer-based dataset together with benchmarks on both established and new datasets.
Such transformer-based datasets that reflect emerging retrieval-augmented generation (RAG) workloads are currently missing in the ANNS and FANNS literature. We generate our embedding vectors with the stella_en_400M_v5 model (NovaSearch, 2025; Zhang et al., 2024), widely adopted in LLM research (Besta et al., 2024; Kim et al., 2025). Unlike our arxiv-for-fanns dataset, end-to-end RAG benchmarks such as the Massive Text Embedding Benchmark (Muennighoff et al., 2022) contain text passages rather than vectors, making them unsuitable for direct evaluation of FANNS methods.
In the broader ANNS field, surveys by Echihabi et al. (Echihabi et al., 2021) and Han et al. (Han et al., 2023) provide comprehensive overviews, and large-scale benchmarks (Aumüller et al., 2020; Aumueller et al., 2025; Wang et al., 2021c) compare methods across datasets. To the best of our knowledge, these do not cover transformer-based embeddings; since our dataset also contains the ground truth for unfiltered ANNS, it could be used to extend these prior works.
@@ 7. Conclusion
The rapid progress of embedding models for text, image, audio, and video has created strong demand for fast and accurate FANNS methods. Recent work has proposed diverse approaches supporting EM, R, and EMIS filtering. To structure this evolving field, we present a taxonomy classifying methods by filtering approach, indexing technique, and filter type, and we survey FANNS methods accordingly. By analyzing how FANNS methods are evaluated in literature, we identify a key limitation: the lack of open datasets with transformer-based embeddings and real-world attributes. Analyzing embeddings from images, audio, video, and text, we find that the embedding model strongly shapes vector characteristics and hence FANNS performance, which motivates the need for transformer-based datasets.
To address this gap, we release the arxiv-for-fanns dataset, which contains over 2.7 million 4096-dimensional vectors with 11 attributes and reflects realistic transformer-based workloads. Using this dataset, we perform an in-depth benchmarking study of FANNS methods, analyzing their performance across different filter types, dataset scales, numbers of retrieved neighbors, and query selectivities. By distilling our results into eight key observations, we provide practical guidance for selecting and configuring FANNS methods, and by publishing our dataset, we establish a benchmark that helps steer future research towards more efficient FANNS methods for state-of-the-art, transformer-based embedding vectors.